OpenStudy (anonymous):

Given the following functions, find g ( f (9) ). f(x)= (1)/(x-10) g(x)=(1)/(x)

OpenStudy (anonymous):

If I plug 9 into f(x) I get(1)/(9-10) f(x)= (1)/(-1) ?

OpenStudy (anonymous):

but then wouldn't that be 0 ?

OpenStudy (math&ing001):

1/(-1)=-1 So you just plug that value in g(x). The final value would be g(-1)